# STeP: A General and Scalable Framework for Solving Video Inverse Problems with Spatiotemporal Diffusion Priors

### v1

@@ 4.1 Baselines & Metrics
Methods PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) d-PSNR ($\uparrow$) d-SSIM ($\uparrow$) FVD ($\downarrow$) Data Misfit
STeP (ours) 38.85 (1.50) 0.96 (0.01) 0.089 (0.019) 45.61 (2.45) 0.98 (0.01) 2153.34 10.31 (0.98)
IDM+BCS 37.51 (1.24) 0.95 (0.01) 0.095 (0.018) 42.92 (1.82) 0.96 (0.01) 2683.83 10.63 (0.94)
IDM 37.70 (0.99) 0.95 (0.01) 0.095 (0.018) 42.73 (1.65) 0.96 (0.01) 2789.80 10.61 (0.94)
Figure 5: Visual results of dynamic MRI. Left: We visualize the complex amplitude of a few selected frames of the ground truth and generated video reconstructions, including optical flows and zoom-ins. Each image is scaled according to the 95-th percentile calculated from the video dataset (i.e. around $95\%$ values are within the range $[0,1]$). STeP provides more accurate reconstructions with better main structure (the valve in the pink box) and fine-grained details (the light-color tissue in the green box). Right: Similarly to the black hole results, we show the temporal dynamics by visualizing the averaged delta between consecutive frames over 4 frames and 12 frames and the average optical flow estimated with TV-L1 optical flow [78]. STeP matches most of the motion patterns in the ground truth.
4.2 Black hole video reconstruction
Problem setup.
The goal is to reconstruct a video ${\mathbf{x}}_{0}\in\mathbb{R}^{n_{f}\times n_{h}\times n_{w}}$ of a rapidly moving black hole. Each measurement, or visibility, is given by correlating the measurements from a pair of telescopes to sample a particular spatial Fourier frequency of the source with very long baseline interferometry (VLBI) [62, 80]. Mathematically, the measured visibility given by the telescope pair $\{a,b\}$ for the $j$-th frame is:
${\mathbf{V}}_{\{a,b\}}^{[j]}=g_{a}^{[j]}g_{b}^{[j]}e^{-i\left(\phi_{a}^{[j]}-\phi_{b}^{[j]}\right)}{\mathbf{I}}_{\{a,b\}}^{[j]}({\mathbf{x}}_{0})+{\mathbf{n}}_{\{a,b\}}^{[j]},$ (5)
where ${\mathbf{I}}_{\{a,b\}}^{[j]}({\mathbf{x}}_{0})\in\mathbb{C}$ is the corresponding ideal visibility. Notably, ${\mathbf{V}}_{\{a,b\}}^{[j]}$ is a corrupted version of ${\mathbf{I}}_{\{a,b\}}^{[j]}({\mathbf{x}}_{0})$ that experiences Gaussian thermal noise ${\mathbf{n}}_{\{a,b\}}^{[j]}$ as well as telescope-dependent amplitude errors $g_{a}^{[j]}$, $g_{b}^{[j]}$ and phase errors $\phi_{a}^{[j]}$, $\phi_{b}^{[j]}$ [20]. To mitigate the impact of these amplitude and phase errors, closure quantities are derived and used to constrain inference [4]. Specifically, closure phases and log closure amplitudes are considered and can be written as:
${\mathbf{y}}_{\text{cp},\{a,b,c\}}^{[j]}$ $=\angle\left({\mathbf{V}}_{\{a,b\}}^{[j]}{\mathbf{V}}_{\{b,c\}}^{[j]}{\mathbf{V}}_{\{a,c\}}^{[j]}\right)\in\mathbb{R},$ (6)
${\mathbf{y}}_{\text{logca},\{a,b,c,d\}}^{[j]}$ $=\log\left(\frac{\left|{\mathbf{V}}_{\{a,b\}}^{[j]}\right|\left|{\mathbf{V}}_{\{c,d\}}^{[j]}\right|}{\left|{\mathbf{V}}_{\{a,c\}}^{[j]}\right|\left|{\mathbf{V}}_{\{b,d\}}^{[j]}\right|}\right)\in\mathbb{R}.$ (7)
Here, ($\angle$($\cdot$)∠ ( ⋅ ) and ||$\cdot$|| ⋅ | denote the complex angle and amplitude. The overall forward model is a combination of the two groups of closure quantities and an additional flux constraint (see Appendix B for more details). The likelihood function $p({\mathbf{y}}\mid{\mathbf{x}}_{0})$ is given by Eq. 16.
Dataset & spatiotemporal prior.
Measurements are simulated under observational conditions similar to those of the real data currently available for black hole video reconstruction. Namely, the Event Horizon Telescope (EHT) array observed the black hole Sagittarius A* over the course of a night in 2017, with approximately 100 minutes of that observation used for video reconstruction in [22]. A dataset of simulated black hole videos is compiled to match the expected dynamics of Sagittarius A* over this timescale. Specifically, we consider general relativistic magnetohydrodynamic (GRMHD) simulations [70] of the Sagittarius A* black hole under different black hole model assumptions and viewing conditions. The entire dataset contains 648 black hole videos, each with 1000 frames at 400$\times$400 spatial resolution. We then downsampled to 64 frames at 256$\times$256 spatial resolution, so $n_{f}=64$. We adopt a 64×64$\times$64 × (8×8$\times$8 × for both height and width) compression encoder and decoder. Details of the training hyperparameters are shown in Tab. 4.
Results.
We show the quantitative results in Tab. 1, and qualitative comparisons in Fig. 4. More results are provided in Appendix E. Quantitative evaluation shows that STeP significantly outperforms the baseline in temporal consistency. Visualizing the averaged delta frames reveals that our generated videos exhibit motion patterns closely matching the ground truth. In contrast, IDM+BCS produces more static motion due to its implicit static temporal prior assumption, while IDM lacks temporal consistency, leading to high incoherence across individual frames. This inconsistency arises because measurements are extremely sparse per frame. By leveraging a spatiotemporal prior, STeP jointly fits measurements across the entire video.
Multi-modal posterior analysis.
As discussed earlier, black hole imaging is a non-convex and highly ill-posed problem with extremely sparse measurements. Our experimental results in Fig. 1 indicate that its posterior distribution can be multi-modal. This implies that generated samples may align with distinct modes that, while differing significantly from the true videos, still fit the measurement data well. These findings demonstrate that STeP can generate diverse yet equally plausible videos, which is desirable for scientific discovery and uncertainty quantification.
4.3 Dynamic MRI
Problem setup.
We consider a dynamic MRI reconstruction problem in cardiac imaging, where the objective is to recover a video ${\mathbf{x}}_{0}\in\mathbb{C}^{n_{f}\times n_{h}\times n_{w}}$ of the heart from the subsampled Fourier space (a.k.a $k$-space) measurements ${\mathbf{y}}$. Mathematically, this can be formulated as
${\mathbf{y}}^{[j]}={\mathbf{m}}^{[j]}\odot\mathcal{F}\left({\mathbf{x}}_{0}^{[j]}\right)+{\mathbf{n}}^{[j]}\in\mathbb{C}^{n}\quad\text{for}j=1,...,n_{f},$
where ${\mathbf{m}}^{[j]}\in\{0,1\}^{n_{h}\times n_{w}}$ is the subsampling mask for the $j$-th frame, $\odot$ denotes element-wise multiplication, $\mathcal{F}$ is the Fourier transform, and ${\mathbf{n}}^{[j]}$ is the measurement noise. In our experiments, we used subsampling masks with an equi-spaced pattern (similar to those visualized in [65]) of both 6×6$\times$6 × acceleration with 24 auto-calibration signal (ACS) lines (Tab. 2) and 8×8$\times$8 × acceleration with 12 ACS lines (Fig. 6). For dynamic MRI, we use the Gaussian likelihood function: $\log p({\mathbf{y}}|{\mathbf{x}}_{0})\propto-\|\mathcal{A}({\mathbf{x}}_{0})-{\mathbf{y}}\|^{2}_{2}$.
Dataset & spatiotemporal prior.
We use the publicly available cardiac cine dataset from the CMRxRecon Challenge 2023 [65]. The entire dataset contains 3,324 cardiac MRI sequences with fully sampled and ECG-triggered $k$-space data from 300 patients, including various canonical views in cardiac imaging. The cardiac cycle was segmented into 12 temporal states, making each scan a 2D video of 12 frames, i.e. $n_{f}=12$. Given the fully sampled $k$-space data, we obtain the target videos by taking the inverse Fourier transform and resize all videos to the same spatial dimension of 192$\times$192. The measurements were generated by retrospectively applying the subsampling mask $\{{\mathbf{m}}^{[j]}\}_{j=1}^{n_{f}}$ to the fully sampled $k$-space data. We adopt a 16×16$\times$16 × (4×4$\times$4 × for both height and width) compression encoder and decoder. The detailed training hyperparameters are shown in Tab. 4.
Results.
We present the quantitative results for dynamic MRI in Tab. 2, with qualitative comparisons shown in Fig. 5. Additional results are provided in Appendix E. Unlike the ill-posed and non-convex nature of black hole imaging, dynamic MRI is a linear inverse problem focused on recovering fine-grained details. To highlight this, we zoom in on relevant structures in both video frames and averaged delta frames visualizations. Quantitative evaluation further demonstrates that STeP significantly outperforms the baseline in temporal consistency, which also enhances spatial alignment in the generated videos. More results are shown in Appendix E.
4.4 Effectiveness of Image-Video Joint Training
@@
To better understand the impact of the spatiotemporal prior on solving inverse problems, we evaluate results using various checkpoints of the spatiotemporal UNet, each representing a prior fine-tuned for a different number of epochs. We assess performance using PSNR (blue curve), d-PSNR (red curve), and a data-fitting metric (green curve), as shown in Fig. 6 with a shared horizontal axis indicating the fine-tuning epochs. Since the spatiotemporal UNet is initialized from a pretrained image diffusion model, these curves reveal the gradual enhancement as increasingly stronger spatiotemporal priors are incorporated. The results indicate that temporal consistency and spatial consistency improve in a steady, synchronized manner as the prior undergoes further fine-tuning, evidenced by the close alignment of the blue and red curves. Furthermore, a better spatiotemporal prior enhances data fitting, as shown by the downward trend of the green curve.
@@
Figure 6: Consistent improvement in image-video joint fine-tuning. We evaluate intermediate checkpoints for solving the inverse problems of (a) black hole imaging and (b) dynamic MRI with 8$\times$ acceleration. Both spatial quality (measured by PSNR) and temporal consistency (measured by d-PSNR) show steady improvement. For black hole imaging, the average chi-square metric is computed as the mean of $\tilde{\bm{\chi}}_{\text{cp}}^{2}$ and $\tilde{\bm{\chi}}_{\text{logca}}^{2}$.
@@ 5 Conclusion
We introduced STeP, a general framework for solving video inverse problems (VIPs) with a spatiotemporal diffusion prior. We demonstrated that it is possible to efficiently train a diffusion prior for videos, even with limited video data, enabling seamless integration into an existing PnP method for video inversion. By capturing complex temporal structure in the diffusion prior, our approach eliminates the need for temporal heuristics and enables the recovery of intricate temporal dynamics that resemble those in the training videos. We applied our method to two challenging scientific VIPs—black hole imaging and dynamic MRI—where it outperformed existing approaches in recovering both fine-grained spatial details and underlying temporal relationships. These results highlight that, with our proposed strategy, a diffusion video prior can be leveraged in a straightforward manner to tackle complex video inverse problems.
@@ Appendix A Detailed implementation of STeP
Algorithm 1 STeP: a general and scalable framework for solving video inverse problems with SpatioTemporal Prior
@@
The probability flow ordinary differential equation (PF-ODE) [31] of the diffusion model, given by Eq. 8, governs the continuous increase or reduction of noise in the image when moving forward or backward in time. Here, $\dot{\sigma}_{t}$ denotes the time derivative of $\sigma_{t}$, and $\nabla_{{\mathbf{z}}_{t}}\log p({\mathbf{z}}_{t};\sigma_{t})$ represents the time-dependent score function [51, 54].
@@
Our goal is to solve the probability flow ODE (PF-ODE), as defined in Eq. 8, backward from $t=t_{i}$ to $t=0$, given the intermediate state ${\mathbf{z}}_{t_{i}}$ and the pretrained latent score function ${\mathbf{s}}_{\bm{\theta}}({\mathbf{z}};\sigma)\approx\nabla_{\mathbf{z}}\log p({\mathbf{z}};\sigma)$. Any ODE solver, such as Euler’s method or the fourth-order Runge-Kutta method (RK4) [9], can be used to solve this problem. Following previous conventions [81], we adopt a few-step Euler method for solving it efficiently.
@@
This approximation holds for $t_{\min}\approx 0$, assuming that ${\mathbf{z}}_{0}$ lies close to the clean latent manifold [52]. To improve both convergence speed and approximation accuracy, the MCMC samplers are initialized with the solutions obtained from the previous PF-ODE step, leveraging its outputs as a warm start.
@@
According to Proposition 1 in [81], one can obtain a sample ${\mathbf{z}}_{t_{i-1}}\sim p({\mathbf{z}}_{t_{i-1}}\mid{\mathbf{y}})$ by simply adding Gaussian noise from a sample $\hat{\mathbf{z}}_{0}\sim p({\mathbf{z}}_{0}\mid{\mathbf{z}}_{t_{i}},{\mathbf{y}})$, given ${\mathbf{z}}_{t_{i}}\sim p({\mathbf{z}}_{t_{i}}\mid{\mathbf{y}})$ from last step. Thus we solve the target posterior sampling by gradually sampling from the time-marginal posterior of diffusion trajectory. The full parameters STeP is summarized in Tab. 3. The HMC-related parameters are searched on a leave out validation dataset consisting of 3 videos that are different from the testing videos.
Table 3: Hyper-parameters of STeP for black hole imaging and dynamic MRI. We provide and group the hyper-parameters of Algorithm 1.
Hyper-parameters Black hole imaging Dynamic MRI

### v2

@@ 4.3 Metrics
Tasks Methods PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) d-PSNR ($\uparrow$) d-SSIM ($\uparrow$) FVD ($\downarrow$) Data Misfit ($\downarrow$)
Black hole BIS kwon2025solving 23.79 (1.41) 0.718 (0.047) 0.179 (0.031) 29.26 (1.51) 0.938 (0.015) 1429.42 1.719 (1.277)
BCS kwon2025solving 27.66 (2.04) 0.816 (0.053) 0.124 (0.040) 41.71 (1.99) 0.979 (0.008) 564.43 1.426 (0.784)
Bilinear chang2024how 26.11 (2.05) 0.718 (0.067) 0.151 (0.044) 33.14 (2.33) 0.958 (0.013) 1335.95 1.384 (0.742)
Bicubic chang2024how 25.68 (1.88) 0.730 (0.058) 0.163 (0.041) 31.70 (1.88) 0.952 (0.015) 1521.14 1.736 (1.259)
Nearest chang2024how 25.29 (1.80) 0.754 (0.059) 0.164 (0.042) 30.76 (1.75) 0.943 (0.017) 1171.07 1.691 (1.020)
$\int$-noise chang2024how ; deng2025infiniteresolution 24.90 (1.52) 0.745 (0.058) 0.163 (0.034) 31.87 (1.93) 0.945 (0.017) 1253.81 1.655 (1.020)
GP-Warp daras2024warped 23.98 (1.28) 0.721 (0.043) 0.176 (0.029) 29.21 (1.24) 0.938 (0.014) 1395.15 1.721 (1.225)
STeP (video only) 28.71 (1.81) 0.802 (0.079) 0.120 (0.041) 41.41 (2.39) 0.975 (0.011) 238.36 1.124 (0.136)
STeP (image-video joint) 30.28 (2.71) 0.865 (0.063) 0.095 (0.039) 42.09 (2.63) 0.976 (0.011) 170.67 1.114 (0.154)
MRI (8$\times$) BIS kwon2025solving 35.04 (0.76) 0.889 (0.016) 0.100 (0.012) 38.91 (1.13) 0.918 (0.012) 82.30 9.203 (0.698)
BCS kwon2025solving 35.43 (0.97) 0.893 (0.016) 0.099 (0.012) 39.91 (1.09) 0.931 (0.011) 95.68 9.225 (0.693)
Bilinear chang2024how 35.30 (0.96) 0.896 (0.016) 0.098 (0.012) 39.87 (1.14) 0.931 (0.011) 87.90 9.157 (0.670)
Bicubic chang2024how 35.31 (0.91) 0.896 (0.016) 0.098 (0.011) 40.11 (1.10) 0.934 (0.010) 108.90 9.189 (0.653)
Nearest chang2024how 34.87 (0.90) 0.895 (0.018) 0.099 (0.013) 40.09 (1.14) 0.933 (0.011) 108.57 9.188 (0.656)
$\int$-noise chang2024how ; deng2025infiniteresolution 35.55 (1.03) 0.892 (0.020) 0.099 (0.014) 39.77 (1.36) 0.929 (0.015) 89.59 9.208 (0.693)
GP-Warp daras2024warped 34.49 (0.65) 0.886 (0.016) 0.102 (0.012) 38.71 (1.14) 0.916 (0.013) 92.19 9.209 (0.659)
STeP (video only) 37.00 (1.46) 0.927 (0.019) 0.086 (0.013) 43.50 (2.70) 0.963 (0.015) 75.27 8.817(0.650)
STeP (image-video joint) 39.38 (1.16) 0.951 (0.009) 0.078 (0.011) 44.86 (1.92) 0.974 (0.006) 78.60 8.753 (0.603)
Figure 4: Visual examples of STeP (bottom left) and baselines for black hole video reconstruction. To facilitate analysis of the reconstructed spatiotemporal structures, we present results in three ways: (1) a single frame to illustrate spatial fidelity, (2) an $x$-$t$ slice depicting temporal evolution of a vertical line to evaluate temporal consistency, and (3) the averaged optical flow visualized using the standard color scheme from teed2020raftrecurrentallpairsfield to assess spatiotemporal coherence jointly. Compared to baselines, STeP exhibits clearer alignment with ground truth videos across all aspects.
Figure 5: Visual examples of STeP (bottom left) and baselines for dynamic MRI. We visualize a representative frame along with two zoomed-in regions for each method to better illustrate spatial fidelity. Benefiting from its robust spatiotemporal prior, STeP provide reconstructions with fewer structural artifacts and temporal fluctuations, as indicated by its averaged optical flow aligning more closely with the ground truth. This demonstrates the effectiveness of our learned spatiotemporal prior in enhancing both spatial and temporal consistency.
4.4 Main results  [^1] For better assessment of video reconstruction quality, we strongly encourage the readers to watch our supplementary video.
STeP reconstructs videos with better spatial and temporal coherence.
We provide quantitative results in Tab. 1 and visual comparisons in Fig. 5 and Fig. 5. STeP outperforms baselines in overall video quality across most metrics and generates video reconstructions with significantly better spatiotemporal coherence. Fig. 5 visualizes an $x$-$t$ slice, representing the temporal evolution (horizontal axis) of a spatial slice (marked by the cyan vertical line). We find that the noise warping baselines fail to constrain temporal consistency, as indicated by the fluctuating ring diameters in the $x$-$t$ slices, while BCS provides almost static reconstructions. In contrast, STeP exhibits closer alignment with the ground truth and improved temporal consistency. This is further illustrated by the averaged optical flow visualization, highlighting that STeP faithfully captures the underlying temporal dynamic. Similar trends can be observed in Fig. 5.
Noise warping is less effective in scientific problems.
We observe that the performance of noise-warping strategies in challenging scientific VIPs does not correlate with noise-warping accuracy. For example, although the $\int$-noise approach has demonstrated superior noise-warping capability given optical flow (as shown in chang2024how ), it underperforms simpler strategies such as Bilinear and Bicubic interpolation in challenging VIPs. This discrepancy arises primarily due to: (1) inaccuracies in the optical flow, and (2) the inherent difficulty of effectively manipulating noise in latent space through pixel space optical-flow-guided warping. Consequently, methods relying on precise optical flow and carefully designed warping strategies struggle with tasks requiring high temporal consistency, such as black hole video reconstruction. In contrast, our approach employs a data-driven spatiotemporal prior learned from training data, effectively overcoming these limitations and enabling broader applicability to scientific tasks with significant temporal uncertainty.
Joint fine-tuning benefits both spatial and temporal consistency.
As shown in Tab. 1, STeP with an image-video jointly fine-tuned spatiotemporal prior outperforms both IDM-based baselines and STeP trained video data only. To further illustrate these improvements, we provide a detailed visual comparison in Fig. 7. In Fig. 7 (a), we analyze the temporal dynamics using averaged delta frames, which visualize the averaged differences between consecutive frames over a window. We observe that delta frames from BCS remain largely unchanged as the temporal window expands, indicating limited temporal variation. In contrast, STeP (joint) effectively captures more complex and coherent temporal dynamics. In Fig. 7 (b), we compare STeP (joint) and STeP (video only) with a focus on spatial fidelity. Although their reconstructions have similar average optical flow, STeP (joint) have fewer artifacts and hallucinations as pointed out by the arrows. Such an improvement highlights the value of using image data in both the pre-training and fine-tuning stages.
STeP provides diverse and equally plausible solutions.
Due to the highly ill-posed and extremely sparse measurements in black hole video reconstruction, STeP can produce multiple semantically diverse reconstructions, each visually plausible and consistent with measurement data. Since the true posterior distribution is unknown, directly quantifying mode coverage is infeasible. Instead, we draw 100 i.i.d. samples and cluster them based on spatial appearance and temporal dynamics. As illustrated in Fig. 1, we identify three distinct modes with nearly identical data fidelity (see Tab. 5 exact data misfit values). One recovered mode aligns closely with the ground truth, while the others differ in rotation direction or spatial structure. This demonstrates that our method not only accurately reconstructs the underlying ground truth but also discovers additional plausible solutions.
4.5 Ablation on image-video joint finetuning
@@
To show the effectiveness of the proposed image-video joint finetuning technique, we show the quantitative results of using checkpoints of the spatiotemporal UNet that were fine-tuned for different numbers of epochs. We assess performance using PSNR (blue curve), d-PSNR (red curve), and a data-fitting metric (green curve), as shown in Fig. 7. Since the spatiotemporal UNet is initialized from a pretrained image diffusion model, these curves indicate steady improvement in spatiotemporal consistency and data fitting as the prior is fine-tuned.
Figure 6: Detailed comparison on black hole video reconstruction and dynamic MRI. Left: We compare STeP (joint) and the BCS baseline kwon2025solving by visualizing the averaged delta frames (difference images) over an expanding window. The delta frames given by STeP (joint) better align with the ground truth, indicating better temporal consistency. Right: We also compare the spatial fidelity between STeP (joint) and its variant STeP (video-only). Trained on both images and videos, STeP (joint) provide reconstructions with less spatial hallucinations compared to STeP (video-only).
@@
Figure 7: Consistent improvement in image-video joint fine-tuning. We evaluate intermediate checkpoints of (a) black hole video reconstruction and (b) dynamic MRI (8$\times$). Spatial similarity (measured by PSNR), temporal consistency (measured by d-PSNR) and measurement data fit (measured by data misfit) show steady improvement.
@@ 5 Conclusion
We introduced STeP, a framework designed for scientific video inverse problems (VIPs) that integrates a spatiotemporal diffusion prior into a PnPDP solver. By explicitly modeling spatiotemporal structures, our method eliminates the reliance on temporal heuristics and recovers coherent videos from sparse measurements. We evaluated STeP on two challenging scientific VIPs: black hole video reconstruction and dynamic MRI. STeP consistently outperformed existing methods, achieving better spatial fidelity and more accurate temporal consistency. Our framework demonstrates a general and scalable approach to scientific VIPs by combining sophisticated spatiotemporal diffusion priors with state-of-the-art PnPDP methods.
@@ Appendix A Detailed implementation of STeP
Algorithm 1 STeP: a framework for solving video inverse problems with SpatioTemporal Prior
@@
The probability flow ordinary differential equation (PF-ODE) (karras2022elucidating, ) of the diffusion model, given by Equation (4), governs the continuous increase or reduction of noise in the image when moving forward or backward in time. Here, $\dot{\sigma}_{t}$ denotes the time derivative of $\sigma_{t}$, and $\nabla_{{\mathbf{z}}_{t}}\log p({\mathbf{z}}_{t};\sigma_{t})$ represents the time-dependent score function (song2019generative, ; song2021scorebased, ).
@@
Our goal is to solve the probability flow ODE (PF-ODE), as defined in Equation (4), backward from $t=t_{i}$ to $t=0$, given the intermediate state ${\mathbf{z}}_{t_{i}}$ and the pretrained latent score function ${\mathbf{s}}_{\bm{\theta}}({\mathbf{z}};\sigma)\approx\nabla_{\mathbf{z}}\log p({\mathbf{z}};\sigma)$. Any ODE solver, such as Euler’s method or the fourth-order Runge-Kutta method (RK4) (butcher2008numerical, ), can be used to solve this problem. Following previous conventions (zhang2024improving, ), we adopt a few-step Euler method for solving it efficiently.
@@
This approximation holds for $t_{\min}\approx 0$, assuming that ${\mathbf{z}}_{0}$ lies close to the clean latent manifold (song2020generativemodelingestimatinggradients, ). To improve both convergence speed and approximation accuracy, the MCMC samplers are initialized with the solutions obtained from the previous PF-ODE step, leveraging its outputs as a warm start.
@@
According to Proposition 1 in (zhang2024improving, ), one can obtain a sample ${\mathbf{z}}_{t_{i-1}}\sim p({\mathbf{z}}_{t_{i-1}}\mid{\mathbf{y}})$ by simply adding Gaussian noise from a sample $\hat{\mathbf{z}}_{0}\sim p({\mathbf{z}}_{0}\mid{\mathbf{z}}_{t_{i}},{\mathbf{y}})$, given ${\mathbf{z}}_{t_{i}}\sim p({\mathbf{z}}_{t_{i}}\mid{\mathbf{y}})$ from last step. Thus we solve the target posterior sampling by gradually sampling from the time-marginal posterior of diffusion trajectory. The full hyperparameters and running cost STeP is summarized in Tab. 2. The HMC-related parameters are searched on a leave out validation dataset consisting of 3 videos that are different from the testing videos.
Table 2: Summary of STeP for black hole video reconstruction and dynamic MRI. We provide and group the hyperparameters of Algorithm 1. HMC related parameters are tuned on $3$ leave-out validation videos. The run time and memory are tested using 1 NVIDIA A 100-SCM4-80GB GPU.
Black hole video reconstruction Dynamic MRI
@@
Inference Related
decoder NFE $N_{\text{dec}}$ 1500 1060
diffusion model NFE $N_{\text{dm}}$ 500 400
time (s) per sample 645 332
memory (GB) 48 21